OpenStudy (anonymous):

Melinda has $22 and plans to save $10 each week. George has $10 and plans to save $13 each week. In how many weeks will they have the same amount saved?

OpenStudy (danjs):

initial amount + (#weeks)*(money saved per week) should be equal for both

OpenStudy (anonymous):

sooo.... what do u thank the answer is

OpenStudy (danjs):

for example, melinda 22 + 10*(weeks)

OpenStudy (anonymous):

32

OpenStudy (danjs):

george 10 + 13*(weeks) set both equal, and solve for (weeks)

OpenStudy (anonymous):

62

OpenStudy (danjs):

what work did you do to arrive there? or just guessing?

OpenStudy (anonymous):

They need 4 weeks solve for n 22+10n =10+13n 22+10x4=10+13x4 22+40 = 10+52 62= 62

OpenStudy (danjs):

RIght, 4 weeks, 22+10n =10+13n n=4 kina hard to read what you did though

OpenStudy (anonymous):

cool thank u

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

did u get my message

OpenStudy (anonymous):

http://openstudy.com/users/DanJS

OpenStudy (anonymous):

can u please help me is the answer 4 or 62